OpenStudy (anonymous):

solve for x (3/k)=(1/x)-(2/y)

OpenStudy (mertsj):

\[\frac{3}{k}=\frac{1}{x}-\frac{2}{y}\]

OpenStudy (mertsj):

Add 2/y to both sides

OpenStudy (mertsj):

\[\frac{3}{k}+\frac{2}{y}=\frac{1}{x}\]

OpenStudy (mertsj):

Add the two fractions on the left side

OpenStudy (mertsj):

\[\frac{3y+2k}{ky}=\frac{1}{x}\]

OpenStudy (mertsj):

Take reciprocals of each side.

OpenStudy (mertsj):

\[\frac{ky}{3y+2k}=x\]

OpenStudy (anonymous):

thank you very much...can you help me with one more problem?

OpenStudy (mertsj):

post it

OpenStudy (anonymous):

(5x/2x-3)+4 [add and simplify]

OpenStudy (mertsj):

\[\frac{5x}{2x-3}+4=\frac{5x+4(2x-3)}{2x-3}=\frac{5x+8x-12}{2x-3}=\]

OpenStudy (mertsj):

\[\frac{13x-12}{2x-3}\]